OpenStudy (anonymous):

• Use the recursive formula to make a table of values for 1 ≤ n ≤ 5. Show your calculations. Explain what your table means.

OpenStudy (anonymous):

Here is the starting of the question..... Bill is able to save $35/week after working part-time and paying his expenses. These two formulas show his weekly savings: • f(1) = 35, f(n) = f(1) + f(n-1) for n > 1 • f(n) = 35n

OpenStudy (mathmale):

If f(1) (the initial value of the function) is $35, and each new value of f(n) equals the 1st value ($35) plus the previous value, for n = {2, 3, 4, 5}, the next function value will be f(2)=f(1)+f(2-1) = f(1) + f(1) = ???

OpenStudy (anonymous):

It would equal f(2) right @mathmale and thanks for all the help

OpenStudy (mathmale):

Actually, f(2) would be f(1) + f(1) = $35 + $35.